Linear Systems Problem Solving
3. One equation of a system is 2x-y=5. Write a second equation so that the system has
a) no solution
b) infinitely many solutions
c) one solution, (3, 1). pleaseeeee
Thanks a lotttt

Answer by user_dude2008(1862)

a)
Just change the 5 to something else. So other equation is 2x-y=7

b)
Multiply all numbers by 2. So equation is 4x-2y=10

c)

Make up 2 numbers for left side. So I chose 4 and 5. Left side is now 4x+5y

Plug in (3,1) -----> 4(3)+5(1)=12+5=17

Equation is 4x+5y=17